OpenStudy (anonymous):

(8/x+1) - (3/2) = (8/3x+3)

OpenStudy (anonymous):

x=32/21

OpenStudy (anonymous):

incorrect?

OpenStudy (anonymous):

yep i know why too, just a second let me rework it

OpenStudy (anonymous):

no, i think its right is this the equation?

OpenStudy (anonymous):

Check out that attachment.. thats the problem..

OpenStudy (anonymous):

i appreciate you helping me a lot!!! this stuff is due tonight. and trying to get it done.

OpenStudy (anonymous):

oh yeah, thats different, just a second lol

OpenStudy (anonymous):

x=23/9

OpenStudy (anonymous):

You are a life saver!

OpenStudy (anonymous):

Try this one! -3y + 5 (6y+9) = -1 -5 (4-y)

OpenStudy (anonymous):

that is another one i'm stuck on. :/ i tried 33, -33

OpenStudy (anonymous):

what do i need 33,-33 for?

OpenStudy (anonymous):

Here is this one:

OpenStudy (anonymous):

check to make sure I set the problem up and its the same

OpenStudy (anonymous):

I tried 33, and -33 for the answers but its not correct..